OpenStudy (bananas1234):

Factor completely. 2x^2y + 12xy +18y

OpenStudy (zzr0ck3r):

hint: \(2x^2y+12xy+18y = 2x^2y+6xy+6xy+18y\)

OpenStudy (zzr0ck3r):

Hint: 2 \[2x^2y+12xy+18y = [2x^2y+6xy]+[6xy+18y] = 2xy(x+3)+6y(x+3)\]

OpenStudy (zzr0ck3r):

Does this help?

OpenStudy (bananas1234):

i think so.

OpenStudy (anonymous):

@zzr0ck3r I think you made it alot more complicated than it really is.. All you have to do, @Bananas1234, is follow PEMDAS

OpenStudy (zzr0ck3r):

I am not sure what you are talking about....

OpenStudy (anonymous):

(aka, 1. Parenthesis, 2. Exponents, 3. Multiplying 4. Division 5. Addition 6. Subraction

OpenStudy (zzr0ck3r):

how does that help us factor ?

OpenStudy (bananas1234):

true, this is different than normal problem solving.

OpenStudy (bananas1234):

if i just used pemdas i probably would not end up with the correct answer because i need to factor. Thanks anyways @AliceCullen :) help is always appreciated

OpenStudy (anonymous):

Oh.. Crap Sorry!! @Bananas1234

OpenStudy (bananas1234):

no problem! :D thanks anyway

OpenStudy (bananas1234):

so do i need to simplify? @zzr0ck3r

jimthompson5910 (jim_thompson5910):

do you see how \[\Large 2xy(x+3)+6y(x+3)\] was formed?

OpenStudy (bananas1234):

yes, i think so

jimthompson5910 (jim_thompson5910):